IF x^4+1/x^4=47 , find the value of (x-1/x)^2. PLEASE ANSWER THIS QUESTION ASAP

Grade:9

1 Answers

Rashi Mishra
3 years ago
Adding 2 on both sides
x4+1/x4+2=47+2
x4+1/x4+2(x4)(1/x4)=49 [a2+b2+2ab=(a+b)2]
(x2+1/x2)2=72
x2+1/x2=7
subtracting 2 on both sides
x2+1/x2-2=7-2
x2+1/x2-2(x)(1/x)=5 [a2+b2-2ab=(a-b)2]
(x-1/x)2=5